OpenStudy (anonymous):

1/4 divided by 3/8 = ?

OpenStudy (anonymous):

1/3 2/3 3/4 3/8

OpenStudy (alexandervonhumboldt2):

\[\frac{ 1 }{ 4 }/\frac{ 3 }{ 8 }=\frac{ 1 }{ 4 }*\frac{ 8 }{ 3 }=\frac{ 1*8 }{ 4*3 }=simplify\]

OpenStudy (anonymous):

252

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

can u explain on how to get the answer

OpenStudy (alexandervonhumboldt2):

\[\frac{ 1 }{ 4 }/\frac{ 3 }{ 8 }=\frac{ 1 }{ 4 }*\frac{ 8 }{ 3 }=\frac{ 1*8 }{ 4*3 }=\[\frac{ 8 }{ 12 }=\frac{ 4*2 }{ 4*3 }=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[\Large{\frac{\frac{ 1 }{ 4 }}{\frac{ 3 }{ 8 }}=\frac{\frac{ 1 }{ 4 }\cdot \frac{ 8 }{ 3}}{\frac{ 3 }{ 8 }\cdot \frac{ 8 }{ 3} }=\frac{\frac{ 1 }{ 4 }\cdot \frac{ 8 }{ 3}}{\frac{ \cancel{3} }{ \cancel{8} }\cdot \frac{ \cancel{8} }{ \cancel{3}} }=\frac{\frac{ 1 }{ 4 }\cdot \frac{ 8 }{ 3}}{1 }}=\frac{ 1 }{ 4 }\cdot \frac{ 8 }{ 3}\]